# Extragradient methods with complexity guarantees for hierarchical variational inequalities

### v1

@@ 3.1 Statement of the main results
Theorem 3.1 derives a rate in terms of the feasibility and the optimality gap of the hierarchical VI problem (P). These rates are established relative to arbitrary compact sets $\mathcal{U}_{1},\mathcal{U}_{2}\subset\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$. Such bounds are meaningful because of the boundedness of the sequence $(z^{k})_{k\geq 1}$, a fact we establish with the derivation of (3.6). The constants involved in (3.6) and (3.7) depend on the Lipschitz modulus of the operators ${\mathsf{F}}_{i}$, and diameter-like constants, which can be exhibited thanks to the boundedness of the trajectory. $\Diamond$
Choosing a specific regularization sequence $(\sigma_{k})_{k\geq 1}$ allows us to turn the estimates from Theorem 3.1 into concrete complexity bounds. In detail, we consider the constant step size policy $t_{k}=t$ and regularization sequence
@@
Let the same Assumptions as in Theorem 3.1 be in place. Assume that the regularization sequence $(\sigma_{k})_{k\geq 1}$ is chosen according to (3.9), and $t_{k}=t\leq\frac{1}{\sqrt{8}L_{1}}$ for all $k\geq 1$. Then, we have
@@
$-B_{\mathcal{U}_{1}}\left[\frac{\sup_{z\in\mathcal{U}_{2}}\norm{z^{1}-z}^{2}}{2Kt(\alpha/\rho)}+\frac{aC_{\mathcal{U}_{2}}}{(\alpha/\rho)(1-\delta)(K+b)^{\delta}}\right]^{1/\rho}\overset{(*)}{\leq}\Theta_{\rm Opt}(\bar{z}^{K}|\mathcal{U}_{1}\cap\mathcal{S}_{2})\leq\frac{C_{\mathcal{U}_{1}}}{a(K+b)^{1-\delta}},$ (3.11)
where the inequality ()(*) holds under the $(\alpha,\rho)$-weakly sharpness assumption.
Remark $3.$
Under the regularization sequence (3.9) and if $\mathcal{S}_{2}$ is $(\alpha,\rho)$-weak sharp, a sufficient condition for Assumption 3 to hold is $1>\delta>1-\frac{1}{\rho}$ (cf. Remark 3.1). This agrees with the analysis in the potential case of [boct2025accelerating]. $\Diamond$
@@ 4.1 Energetic bounds
In this section, we develop the necessary energy bounds for the sequence $(z^{k})_{k\geq 1}$ generated by the optimistic extragradient method. Key is the step size condition $8t_{k}^{2}L^{2}_{k}\leq 1$, where $L_{k}\triangleq L_{{\mathsf{F}}_{2}}+\sigma_{k}L_{{\mathsf{F}}_{1}}$ is the combined Lipschitz constant of the operator ${\mathsf{V}}_{k}$. This condition implies $t_{k}\leq\frac{1}{\sqrt{8}}\frac{1}{L_{{\mathsf{F}}_{2}}+\sigma_{k}L_{{\mathsf{F}}_{1}}}\in\frac{1}{\sqrt{8}}\left[\frac{1}{L_{{\mathsf{F}}_{2}}+\sigma_{1}L_{{\mathsf{F}}_{1}}},\frac{1}{L_{{\mathsf{F}}_{2}}}\right].$
@@
Let $(\sigma_{k})_{k\geq 1}$ be non-increasing and $(t_{k})_{k\geq 1}$ be a step size policy. Let also $z\in\mathcal{Z}$ be arbitrary.
  • (i)
    For all $k\geq 1$ we have
    $\begin{split}\frac{1}{2}\norm{z^{k+1}-z}^{2}&\leq\frac{1}{2}\norm{z^{k}-z}^{2}-\frac{1}{2}\norm{z^{k+1/2}-z^{k}}^{2}+\frac{t^{2}_{k}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}\\ &+t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2}),z-z^{k+1/2}\rangle+t_{k}\left(G_{k}(z)-G_{k}(z^{k+1/2})\right).\end{split}$ (4.1)
  • (ii)
    If $8t^{2}_{k}L_{k}^{2}\leq 1$ then for all $k\geq 2$,
    $\begin{split}\frac{1}{2}\norm{z^{k+1}-z}^{2}+\frac{t^{2}_{k}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}&\leq\frac{1}{2}\norm{z^{k}-z}^{2}+\frac{t^{2}_{k-1}}{2}\norm{{\mathsf{V}}_{k-1}(z^{k-3/2})-{\mathsf{V}}_{k-1}(z^{k-1/2})}^{2}\\ &-t_{k}\left(\langle{\mathsf{V}}_{k}(z^{k+1/2}),z^{k+1/2}-z\rangle+G_{k}(z^{k+1/2})-G_{k}(z)\right)\\ &-\frac{1}{4}\norm{z^{k+1/2}-z^{k}}^{2}.\end{split}$ (4.2)
@@
(i) Applying (2.1) to the iterates $z^{k+1/2}$ and $z^{k+1}$, respectively, we obtain
$\langle z^{k+1}-z^{k},z^{k+1}-z\rangle\leq\langle t_{k}{\mathsf{V}}_{k}(z^{k+1/2}),z-z^{k+1}\rangle+t_{k}\left(G_{k}(z)-G_{k}(z^{k+1})\right)$ (4.3)
$\langle z^{k+1/2}-z^{k},z^{k+1/2}-z\rangle\leq\langle t_{k}{\mathsf{V}}_{k}(z^{k-1/2}),z-z^{k+1/2}\rangle+t_{k}\left(G_{k}(z)-G_{k}(z^{k+1/2})\right).$
Using $z=z^{k+1}$ in the second inequality, we have
$\langle z^{k+1/2}-z^{k},z^{k+1/2}-z^{k+1}\rangle\leq\langle t_{k}{\mathsf{V}}_{k}(z^{k-1/2}),z^{k+1}-z^{k+1/2}\rangle+t_{k}\left(G_{k}(z^{k+1})-G_{k}(z^{k+1/2})\right).$ (4.4)
@@
$\norm{z^{k+1}-z}^{2}$ $=\norm{z^{k+1}-z^{k}+z^{k}-z}^{2}$
$=\norm{z^{k}-z}^{2}-\norm{z^{k+1}-z^{k}}^{2}+2\langle z^{k+1}-z^{k},z^{k+1}-z\rangle.$
In the same spirit, we see
$\norm{z^{k+1}-z^{k}}^{2}=\norm{z^{k+1}-z^{k+1/2}}^{2}+\norm{z^{k+1/2}-z^{k}}^{2}+2\langle z^{k+1}-z^{k+1/2},z^{k+1/2}-z^{k}\rangle.$
Combining these two identities with (4.3) and (4.4), we can continue the estimation as
@@
$+2t_{k}\langle{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2}),z^{k+1}-z^{k+1/2}\rangle+2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2}),z-z^{k+1/2}\rangle$
$+2t_{k}\left(G_{k}(z)-G_{k}(z^{k+1/2})\right).$
Applying Young’s inequality $\langle a,b\rangle\leq\frac{c}{2}\norm{a}^{2}+\frac{1}{2c}\norm{b}^{2}$ with $c=t_{k}$, we obtain
$2t_{k}\langle{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2}),z^{k+1}-z^{k+1/2}\rangle\leq t^{2}_{k}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}+\norm{z^{k+1}-z^{k+1/2}}^{2}.$
Hence, we obtain (4.1)
$\frac{1}{2}\norm{z^{k+1}-z}^{2}$ $\leq\frac{1}{2}\norm{z^{k}-z}^{2}-\frac{1}{2}\norm{z^{k+1/2}-z^{k}}^{2}+\frac{t^{2}_{k}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}$
$+t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2}),z-z^{k+1/2}\rangle+t_{k}\left(G_{k}(z)-G_{k}(z^{k+1/2})\right).$
(ii) Observe that for $L_{k}>0$,
$\frac{t_{k}^{2}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}$ $=t^{2}_{k}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}-\frac{t^{2}_{k}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}$
$\leq t_{k}^{2}L_{k}^{2}\norm{z^{k+1/2}-z^{k-1/2}}^{2}-\frac{t^{2}_{k}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}$
$\leq 2t_{k}^{2}L^{2}_{k}\norm{z^{k+1/2}-z^{k}}^{2}+2t_{k}^{2}L^{2}_{k}\norm{z^{k}-z^{k-1/2}}^{2}$
$-\frac{t^{2}_{k}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2},$
@@
$\frac{t^{2}_{k}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}-\frac{1}{2}\norm{z^{k+1/2}-z^{k}}^{2}$ $\leq 2t_{k}^{2}L^{2}_{k}\norm{z^{k}-z^{k-1/2}}^{2}-\frac{t^{2}_{k}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}$
$+\frac{1}{2}\left(4t_{k}^{2}L^{2}_{k}-1\right)\norm{z^{k+1/2}-z^{k}}^{2}$ (4.6)
$\leq 2t_{k}^{2}L^{2}_{k}\norm{z^{k}-z^{k-1/2}}^{2}-\frac{t^{2}_{k}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}$
$-\frac{1}{4}\norm{z^{k+1/2}-z^{k}}^{2},$ (4.7)
where the last inequality is due to the step size choice satisfying $4t_{k}^{2}L^{2}_{k}\leq\frac{1}{2}$. Moreover, using the definition of $z^{k}$ and $z^{k-1/2}$, the non-expansiveness of the proximal mapping (LABEL:eq:nonexpansive), and the step size assumption $8t^{2}_{k}L_{k}^{2}\leq 1$ yields $4t^{2}_{k}t^{2}_{k-1}L^{2}_{k}\leq\frac{t^{2}_{k-1}}{2}$. This results in
$2t_{k}^{2}L_{k}^{2}\norm{z^{k}-z^{k-1/2}}^{2}\leq 2t^{2}_{k}t^{2}_{k-1}L^{2}_{k}\norm{{\mathsf{V}}_{k-1}(z^{k-3/2})-{\mathsf{V}}_{k-1}(z^{k-1/2})}^{2}\leq\frac{t^{2}_{k-1}}{2}\norm{{\mathsf{V}}_{k-1}(z^{k-3/2})-{\mathsf{V}}_{k-1}(z^{k-1/2})}^{2}.$ (4.8)
Combining this inequality with (4.7), plugging the result into (4.1), and rearranging terms, we obtain (4.2)
$\frac{1}{2}\norm{z^{k+1}-z}^{2}+\frac{t^{2}_{k}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}$ $\leq\frac{1}{2}\norm{z^{k}-z}^{2}+\frac{t^{2}_{k-1}}{2}\norm{{\mathsf{V}}_{k-1}(z^{k-3/2})-{\mathsf{V}}_{k-1}(z^{k-1/2})}^{2}$
$-t_{k}\left(\langle{\mathsf{V}}_{k}(z^{k+1/2}),z^{k+1/2}-z\rangle+G_{k}(z^{k+1/2})-G_{k}(z)\right)$
$-\frac{1}{4}\norm{z^{k+1/2}-z^{k}}^{2}.$
$\blacksquare$
To continue with our energetic bounds, we define the data
$D_{k}\triangleq\frac{t^{2}_{k-1}}{2}\norm{{\mathsf{V}}_{k-1}(z^{k-3/2})-{\mathsf{V}}_{k-1}(z^{k-1/2})}^{2}\hskip 17.00024pt\forall k\geq 2.$ (4.11)
This allows us to rewrite the statement in Lemma 4.1(ii) concisely as
$E_{k+1}(z)+D_{k+1}\leq E_{k}(z)+D_{k}-t_{k}\Psi_{k}(z)-\frac{1}{4}\norm{z^{k+1/2}-z^{k}}^{2}\hskip 17.00024pt\forall k\geq 2.$ (4.12)
We now extend the validity of (4.12) to $k\geq 1$.
Lemma 4.2.
Define $D_{1}\triangleq 0$. Then, for all $k\geq 1$ relation (4.12) holds.
Proof.
It is only left to show that the inequality holds for $k=1$. For $k=1$ (4.12) reduces to
$E_{2}(z)+D_{2}\leq E_{1}(z)+D_{1}-t_{1}\Psi_{1}(z)-\frac{1}{4}\norm{z^{3/2}-z^{1}}^{2}.$ (4.13)
Using (4.7) we have that
$\frac{t_{1}^{2}}{2}\norm{{\mathsf{V}}_{1}(z^{1/2})-{\mathsf{V}}_{1}(z^{3/2})}^{2}-\frac{1}{2}\norm{z^{3/2}-z^{1}}^{2}$ $\leq 2t_{1}^{2}L^{2}_{1}\norm{z^{1}-z^{1/2}}^{2}-\frac{t^{2}_{1}}{2}\norm{{\mathsf{V}}_{1}(z^{1/2})-{\mathsf{V}}_{1}(z^{3/2})}^{2}-\frac{1}{4}\norm{z^{3/2}-z^{1}}^{2}$
$=-\frac{t^{2}_{1}}{2}\norm{{\mathsf{V}}_{1}(z^{1/2})-{\mathsf{V}}_{1}(z^{3/2})}^{2}-\frac{1}{4}\norm{z^{3/2}-z^{1}}^{2}$
where the last equality is due to $z^{1/2}=z^{1}$. Plugging this result back into (4.1), we arrive at
$\frac{1}{2}\norm{z^{2}-z}^{2}$ $\leq\frac{1}{2}\norm{z^{1}-z}^{2}-t_{1}\left(\langle{\mathsf{V}}_{1}(z^{3/2}),z^{3/2}-z\rangle+G_{1}(z^{3/2})-G_{1}(z)\right)-\frac{t_{1}^{2}}{2}\norm{{\mathsf{V}}_{1}(z^{1/2})-{\mathsf{V}}_{1}(z^{3/2})}^{2}$
$-\frac{1}{4}\norm{z^{3/2}-z^{1}}^{2}.$
Using the definition of $D_{1}$, we arrive at (4.13). $\blacksquare$
@@
To obtain a bound on the feasibility gap, we start with an Opial-like Lemma which holds under Assumption 3.
@@
Let $z^{*}\in\mathcal{S}_{1}$ arbitrary and $(z^{k},z^{k+1/2})_{k\geq 1}$ be the sequence generated by Algorithm 1. Let also Assumption 3 hold. Define
$W_{k}(z^{*})\triangleq E_{k}(z^{*})+D_{k}\hskip 17.00024pt\forall k\geq 1.$ (4.14)
@@
    $\lim_{k\to\infty}\norm{z^{k+1/2}-z^{k}}=0$;
@@
Pick $z^{\ast}\in\mathcal{S}_{1}\subseteq\mathcal{S}_{2}$ so that there exists $p^{\ast}\in\operatorname{\mathsf{NC}}_{\mathcal{S}_{2}}(z^{\ast})$ such that ${\mathsf{F}}_{1}(z^{\ast})-p^{\ast}\in\partial g_{1}(z^{\ast})$. Using the monotonicity of ${\mathsf{F}}_{1}$ and the convex subgradient inequality for $g_{1}$, we obtain
$\langle{\mathsf{F}}_{1}(z^{k+1/2}),z^{k+1/2}-z^{\ast}\rangle$ $+g_{1}(z^{k+1/2})-g_{1}(z^{\ast})+\langle p^{\ast},z^{k+1/2}-z^{\ast}\rangle$
@@
Combined with the energy inequality (4.12) we conclude
$E_{k+1}(z^{\ast})+D_{k+1}$ $\leq E_{k+1}(z^{\ast})+D_{k+1}+t_{k}\sigma_{k}\left(\langle{\mathsf{F}}_{1}(z^{k+1/2}),z^{k+1/2}-z^{\ast}\rangle+g_{1}(z^{k+1/2})-g_{1}(z^{\ast})+\langle p^{\ast},z^{k+1/2}-z^{\ast}\rangle\right)$
$\leq E_{k}(z^{\ast})+D_{k}-\frac{1}{4}\norm{z^{k+1/2}-z^{k}}^{2}$
$-t_{k}\left(\langle{\mathsf{F}}_{2}(z^{k+1/2}),z^{k+1/2}-z^{\ast}\rangle+g_{2}(z^{k+1/2})-g_{2}(z^{\ast})-\langle\sigma_{k}p^{\ast},z^{k+1/2}-z^{\ast}\rangle\right).$ (4.15)
@@
$-\left(\langle{\mathsf{F}}_{2}(z^{k+1/2}),z^{k+1/2}-z^{\ast}\rangle+g_{2}(z^{k+1/2})-g_{2}(z^{\ast})-\langle\sigma_{k}p^{\ast},z^{k+1/2}-z^{\ast}\rangle\right)$
$=\langle\sigma_{k}p^{\ast},z^{k+1/2}\rangle+\langle{\mathsf{F}}_{2}(z^{k+1/2}),z^{\ast}-z^{k+1/2}\rangle+g_{2}(z^{\ast})-g_{2}(z^{k+1/2})-\langle\sigma_{k}p^{\ast},z^{\ast}\rangle$

### v2

@@ 3.1 Statement of the main results
Theorem 3.1 derives a rate in terms of the feasibility and the optimality gap of the hierarchical HVI problem (P). These rates are established relative to arbitrary compact sets $\mathcal{U}_{1},\mathcal{U}_{2}\subset\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$. Such bounds are meaningful because of the boundedness of the sequence $(z^{k})_{k\geq 1}$, a fact we establish with the derivation of (3.6). The constants involved in (3.6) and (3.7) depend on the Lipschitz modulus of the operators ${\mathsf{F}}_{i}$, and diameter-like constants, which can be exhibited thanks to the boundedness of the trajectory. The stepsize condition is satisfied whenever $t_{k}\leq\frac{1}{4}\frac{1}{L_{{\mathsf{F}}_{2}}+\sigma_{k}L_{{\mathsf{F}}_{1}}}$. This can be achieved by a time-varying stepsize or, since $\sigma_{k}$ is decreasing, by a fixed stepsize, e.g., $t_{k}=t\leq\frac{1}{4}\frac{1}{L_{{\mathsf{F}}_{2}}+\sigma_{1}L_{{\mathsf{F}}_{1}}}$. $\Diamond$
Choosing a specific regularization sequence $(\sigma_{k})_{k\geq 1}$ allows us to turn the estimates from Theorem 3.1 into concrete convergence rates. In detail, we consider the constant step size policy $t_{k}=t$ and regularization sequence
@@
Suppose the same assumptions as in Theorem 3.1 hold. Assume that the regularization sequence $(\sigma_{k})_{k\geq 1}$ is chosen according to (3.9), and $t_{k}=t\leq\frac{1}{4(L_{{\mathsf{F}}_{2}}+\sigma_{1}L_{{\mathsf{F}}_{2}})}$ for all $k\geq 1$. Then, we have
@@
$-B_{\mathcal{U}_{1}}\left[\frac{\sup_{z\in\mathcal{U}_{2}}\norm{z^{1}-z}^{2}}{2Kt(\alpha/\rho)}+\frac{aC_{\mathcal{U}_{2}}}{(\alpha/\rho)(1-\delta)(K+b)^{\delta}}\right]^{1/\rho}$
$\hskip 68.00098pt\overset{(*)}{\leq}\Theta_{\rm Opt}(\bar{z}^{K}|\mathcal{U}_{1}\cap\mathcal{S}_{2})\leq\frac{C_{\mathcal{U}_{1}}}{a(K+b)^{1-\delta}},$ (3.11)
where the inequality ()(*) holds under the $(\alpha,\rho)$-weak sharpness assumption.
Remark 3.4.
Under the regularization sequence (3.9) and $(\alpha,\rho)$-weak sharpness of the lower-level solution set $\mathcal{S}_{2}$, a sufficient condition for Assumption 3 to hold is $1>\delta>1-\frac{1}{\rho}$ (cf. Remark 3.2). This coincides with the analysis for the optimization case in [13, Theorem 2.3]. $\Diamond$
Theorem 3.1 is instrumental to draw conclusions about the asymptotic behavior of the sequences generated by Algorithm 1. However, this statement requires one more (mild) assumption on the data defining problem (P).
Assumption 5.
The non-smooth functions in the definition of (P) satisfy the hierarchical coherence condition $\operatorname{dom}(g_{2})\subseteq\operatorname{dom}(g_{1}).$
Theorem 3.3.
Consider problem (P). Let Assumptions 1-3 and the hierarchic coherence condition Assumption 5 hold. Let $(z^{k})_{k\geq 1}$ and $(z^{k+1/2})_{k\geq 1}$ be generated by Algorithm 1 with $16t_{k}^{2}L_{k}^{2}\leq 1$ for all $k\geq 1$. Then, the following holds true:
  • (i)
    Asymptotic feasibility: All weak accumulation points of $(z^{k+1/2})_{k\geq 1}$ are contained in $\mathcal{S}_{2}$.
  • (ii)
    Ergodic average optimality: If additionally Assumption 4 holds, then the sequence $(\bar{z}^{k})_{k\geq 1}$ converges weakly to an element in $\mathcal{S}_{1}$.
@@ 4.1 Energetic bounds
In this section, we develop the energy bounds for the sequences $(z^{k})_{k\geq 1}$, $(z^{k+1/2})_{k\geq 1}$ generated by the optimistic extragradient method. To simplify the derivations, we introduce the shortcut notation
$D_{k}\triangleq t^{2}_{k-1}\norm{{\mathsf{V}}_{k-1}(z^{k-3/2})-{\mathsf{V}}_{k-1}(z^{k-1/2})}^{2}\hskip 17.00024pt\forall k\geq 2.$ (4.3)
@@
[…]
$\begin{split}E_{k+1}(z)\leq E_{k}(z)-\frac{1}{2}\norm{z^{k+1/2}-z^{k}}^{2}-\frac{1}{4}\norm{z^{k+1}-z^{k+1/2}}^{2}+D_{k+1}-t_{k}\Psi_{k}(z)\end{split}$ (4.4)
If additionally $16t^{2}_{k}L_{k}^{2}\leq 1$ and defining $D_{1}\triangleq 0$, we have, for all $k\geq 1$,
$\begin{split}E_{k+1}(z)+D_{k+1}&\leq E_{k}(z)+D_{k}-t_{k}\Psi_{k}(z)-\frac{1}{4}\norm{z^{k+1/2}-z^{k}}^{2}\\ &-\frac{1}{4}\norm{z^{k+1}-z^{k+1/2}}^{2}.\end{split}$ (4.5)
@@
Proof of (4.4). Applying (2.1) to the iterate $z^{k+1}$, we obtain $\forall z\in\mathcal{Z}$
$\langle z^{k+1}-z^{k},z^{k+1}-z\rangle\leq\langle t_{k}{\mathsf{V}}_{k}(z^{k+1/2}),z-z^{k+1}\rangle+t_{k}\left(G_{k}(z)-G_{k}(z^{k+1})\right).$ (4.6)
Doing the same for the iterate $z^{k+1/2}$, but substituting $z=z^{k+1}$, we get $\forall z\in\mathcal{Z}$
$\begin{split}\langle z^{k+1/2}-z^{k},z^{k+1/2}-z^{k+1}\rangle&\leq\langle t_{k}{\mathsf{V}}_{k}(z^{k-1/2}),z^{k+1}-z^{k+1/2}\rangle\\ &+t_{k}\left(G_{k}(z^{k+1})-G_{k}(z^{k+1/2})\right).\end{split}$ (4.7)
@@
$\norm{z^{k+1}-z}^{2}$ $=\norm{z^{k}-z}^{2}-\norm{z^{k+1}-z^{k}}^{2}+2\langle z^{k+1}-z^{k},z^{k+1}-z\rangle,$ (4.8)
$\norm{z^{k+1}-z^{k}}^{2}$ $=\norm{z^{k+1}-z^{k+1/2}}^{2}+\norm{z^{k+1/2}-z^{k}}^{2}$
$+2\langle z^{k+1}-z^{k+1/2},z^{k+1/2}-z^{k}\rangle.$
Combining these two identities with (4.6) and (4.7), we can continue as
@@
$\norm{z^{k+1}-z}^{2}\leq\norm{z^{k}-z}^{2}-\norm{z^{k+1}-z^{k+1/2}}^{2}-\norm{z^{k+1/2}-z^{k}}^{2}$
$\qquad+2t_{k}\langle{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2}),z^{k+1}-z^{k+1/2}\rangle$
$\qquad+2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2}),z-z^{k+1/2}\rangle+2t_{k}\left(G_{k}(z)-G_{k}(z^{k+1/2})\right).$ (4.10)
Applying Young’s inequality $\langle a,b\rangle\leq\frac{c}{2}\norm{a}^{2}+\frac{1}{2c}\norm{b}^{2}$ with $c=2t_{k}$ and (4.3), we obtain
$2t_{k}\langle{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2}),z^{k+1}-z^{k+1/2}\rangle$
$\leq 2t^{2}_{k}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}+\frac{1}{2}\norm{z^{k+1}-z^{k+1/2}}^{2}$
$=2D_{k+1}+\frac{1}{2}\norm{z^{k+1}-z^{k+1/2}}^{2}.$
Plugging this inequality and definitions (4.1), (4.2) into the previous inequality gives us (4.4).
Proof of (4.5). Observe that by (4.3)
$D_{k+1}$ $=2t^{2}_{k}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}-D_{k+1}\leq 2t_{k}^{2}L_{k}^{2}\norm{z^{k+1/2}-z^{k-1/2}}^{2}-D_{k+1}$
$\leq 4t_{k}^{2}L^{2}_{k}\norm{z^{k+1/2}-z^{k}}^{2}+4t_{k}^{2}L^{2}_{k}\norm{z^{k}-z^{k-1/2}}^{2}-D_{k+1},$
@@
$D_{k+1}-\frac{1}{2}\norm{z^{k+1/2}-z^{k}}^{2}$
$\leq 4t_{k}^{2}L^{2}_{k}\norm{z^{k}-z^{k-1/2}}^{2}-D_{k+1}+\left(4t_{k}^{2}L^{2}_{k}-\frac{1}{2}\right)\norm{z^{k+1/2}-z^{k}}^{2}$ (4.11)
$\leq 4t_{k}^{2}L^{2}_{k}\norm{z^{k}-z^{k-1/2}}^{2}-D_{k+1}-\frac{1}{4}\norm{z^{k+1/2}-z^{k}}^{2},$ (4.12)
where the last inequality is due to the step size choice satisfying $4t_{k}^{2}L^{2}_{k}\leq\frac{1}{4}$. If $k\geq 2$, using the definition of $z^{k}$ and $z^{k-1/2}$, the non-expansiveness of the proximal mapping, and that the step size assumption $16t^{2}_{k}L_{k}^{2}\leq 1$ yields $4t^{2}_{k}L^{2}_{k}\leq 1$, we obtain
$4t_{k}^{2}L_{k}^{2}\norm{z^{k}-z^{k-1/2}}^{2}\leq 4t^{2}_{k}L^{2}_{k}t^{2}_{k-1}\norm{{\mathsf{V}}_{k-1}(z^{k-3/2})-{\mathsf{V}}_{k-1}(z^{k-1/2})}^{2}\leq D_{k}.$ (4.13)
The same inequality holds for $k=1$ since $D_{1}=0$ and $z^{1/2}=z^{1}$. Combining (4.13) with (4.12), plugging the result into (4.4), using the notation (4.2), and rearranging terms, we obtain (4.5). $\blacksquare$
@@
To obtain a bound on the feasibility gap, we start with an Opial-like Lemma.
@@
Let Assumptions 1-3 hold. Let $z^{*}\in\mathcal{S}_{1}$ be arbitrary and let $(z^{k})_{k\geq 1}$ and $(z^{k+1/2})_{k\geq 1}$ be generated by Algorithm 1 with $16t^{2}_{k}L_{k}^{2}\leq 1$. Define $D_{1}=0$ and
$W_{k}(z)\triangleq E_{k}(z)+D_{k}\hskip 17.00024pt\forall k\geq 1.$ (4.14)
@@
    $\lim_{k\to\infty}\norm{z^{k+1/2}-z^{k}}=\lim_{k\to\infty}\norm{z^{k+1}-z^{k+1/2}}=\lim_{k\to\infty}\norm{z^{k+1/2}-z^{k-1/2}}=0$;
@@
  4. (d)
    $\lim_{k\to\infty}\norm{z^{k}-z^{\ast}}$ and $\lim_{k\to\infty}\norm{z^{k+1/2}-z^{*}}$ both exist and are equal.
@@
Pick $z^{\ast}\in\mathcal{S}_{1}\subseteq\mathcal{S}_{2}$ so that there exists $p^{\ast}\in\operatorname{\mathsf{NC}}_{\mathcal{S}_{2}}(z^{\ast})$ s.t. ${\mathsf{F}}_{1}(z^{\ast})-p^{\ast}\in\partial g_{1}(z^{\ast})$. Using the monotonicity of ${\mathsf{F}}_{1}$ and the convex subgradient inequality for $g_{1}$, we obtain
$-H^{({\mathsf{F}}_{1},g_{1})}(z^{\ast},z^{k+1/2})+\langle p^{\ast},z^{k+1/2}-z^{\ast}\rangle$
$=\langle{\mathsf{F}}_{1}(z^{k+1/2}),z^{k+1/2}-z^{\ast}\rangle+g_{1}(z^{k+1/2})-g_{1}(z^{\ast})+\langle p^{\ast},z^{k+1/2}-z^{\ast}\rangle$
@@
Combining with the energy inequality (4.5) and the fact that
$\Psi_{k}(z)=-\sigma_{k}H^{({\mathsf{F}}_{1},g_{1})}(z,z^{k+1/2})-H^{({\mathsf{F}}_{2},g_{2})}(z,z^{k+1/2}),$ (4.15)
we continue
$E_{k+1}(z^{\ast})+D_{k+1}$
$\leq E_{k+1}(z^{\ast})+D_{k+1}+t_{k}\sigma_{k}\left(-H^{({\mathsf{F}}_{1},g_{1})}(z^{\ast},z^{k+1/2})+\langle p^{\ast},z^{k+1/2}-z^{\ast}\rangle\right)$
$\leq E_{k}(z^{\ast})+D_{k}-\frac{1}{4}\norm{z^{k+1/2}-z^{k}}^{2}-\frac{1}{4}\norm{z^{k+1}-z^{k+1/2}}^{2}$
$+t_{k}\left(H^{({\mathsf{F}}_{2},g_{2})}(z^{\ast},z^{k+1/2})+\langle\sigma_{k}p^{\ast},z^{k+1/2}-z^{\ast}\rangle\right).$ (4.16)
@@
$H^{({\mathsf{F}}_{2},g_{2})}(z^{\ast},z^{k+1/2})+\langle\sigma_{k}p^{\ast},z^{k+1/2}-z^{\ast}\rangle$